OpenStudy (anonymous):

HELP?! Thanks to @Miguel_Castillo and @texaschic101 for giving me the answer. But it doesn't fit into my formula. Help?

OpenStudy (anonymous):

What number is 66 2/3% of 81? = 54 My formula (I have to use it) a = 66.67 - ----- 81 100

OpenStudy (whpalmer4):

\[\frac{x\%}{81} = \frac{66.67\%}{100}\] Cross multiply: \[x\%*100 = 81*66.67\%\] \[x\% = \frac{81*66.67\%}{100}\]

OpenStudy (anonymous):

I get 54.0027

OpenStudy (anonymous):

Do I round it?

OpenStudy (e.mccormick):

You have a rounding error already by converting 2/3 to .6667 If you used .6666667 you would see the error get smaller and smaller. In the end, it is 54.

OpenStudy (whpalmer4):

for an exact answer, use 2/3 instead of 66.67%

OpenStudy (anonymous):

Thanks, I tried using decimals but it doesn't work. :/ Thank you to both!